OpenStudy (isabelle99):

Im not sure if the answer i got are right someone help The nth term of a geometric sequence is given by an= 1/4^(4)n-1 . Write the first five terms of this sequence.

OpenStudy (isabelle99):

a1=1/4 a2=1 a3=4 a4=16 I feel like i got these wrong can someone help

OpenStudy (isabelle99):

and a5=64

OpenStudy (sunnnystrong):

Hmmm... Well \[a _{n}=\frac{ 1 }{ 4 }^{4n-1}\] is that correct @Isabelle99 ?

OpenStudy (sunnnystrong):

or is it....... \[a ^{n}=\frac{ 1 }{ 4 }(4)^{n-1}\] because that'd make more sense

OpenStudy (sunnnystrong):

If so: the sequence is \[\frac{ 1 }{ 4 }, 1, 4, 16, 64\] & you're right :D

OpenStudy (isabelle99):

thanks

OpenStudy (sunnnystrong):

NP :D Always happy to help out

OpenStudy (isabelle99):

I didn't even see you previous answer yes thats what i meant to put

OpenStudy (sunnnystrong):

Okay.. well than it's different

OpenStudy (sunnnystrong):

If \[a _{n}=\frac{ 1 }{ 4 }^{4n-1}\] The sequence is: \[\frac{ 1 }{ 4 }^3,\frac{ 1 }{ 4 }^7,\frac{ 1 }{ 4 }^{11},\frac{ 1 }{ 4 }^{15},\frac{ 1 }{ 4 }^{19}...\]

OpenStudy (isabelle99):

no it was the first one you said

OpenStudy (isabelle99):

my bad